Suppose that A and B are disjoint for which P(A) = 0.2 and P(B) = 0.7.
What is the probability that
(a) Either A or B occurs?
(b) A occurs but B does not?
(c) Both A and B occur?

Answer by ikleyn(53937)

.
Suppose that A and B are disjoint for which P(A) = 0.2 and P(B) = 0.7.
What is the probability that

(a) Either A or B occurs?        P(A) + P(B) = 0.2 + 0.7 = 0.9.

(b) A occurs but B does not?     P(A) = 0.2    (! A and B are disjoint ! ! )

(c) Both A and B occur?          This probability is 0 (zero, ZERO).